# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ D.1 Can We Make a DM to Output Any Image With Adversarial Embeddings?
D.2 Comparing Behavioral Differences Between Sets
Our findings from Sec. D.1 undermine our adversarial-based memorization identification. In effect, it may seem that our results regarding NeMo and Wanda (Sec. 3.3) locality in the embedding space (Sec. 4.1). and locality in the model’s weights (Sec. 4.2) become invalid. Indeed, if we are able to trigger generation of any image, then we should not claim that NeMo and Wanda only conceal the memorization instead of fully removing it, and the findings regarding localization would be false, as the obtained adversarial embeddings ${\bm{y}}_{adv}$ yield little information about how the model (and the embedding space) behaves when faced with memorized data.
@@ D.2 Comparing Behavioral Differences Between Sets
In Fig. 5 (right) we show how the SSCD score progresses with the optimization. We note that for verbatim memorization (and generated template memorization) we need only a handful of optimization steps to obtain ${\bm{y}}_{adv}$ that reliably triggers generation of the images. For non-memorized data we reach SSCD above $0.7$ after approximately 500 steps. Notably, we require as much as 200 steps to craft ${\bm{y}}_{adv}$ that reliably produces generated (non-memorized) images.
These results show that the optimization process is indeed different for memorized images than for other images. Building on these findings we allow the optimization procedure to only perform $50$ update steps—a value that guarantees generation of memorized images (if present in the model), while preventing False Positives, i.e., generation of non-memorized images. This constraint ensures methodological correctness of our adversarial-based approach, and proves our results in Secs. 3.3 and 4 are valid.
@@
D.3 Can the Embeddings Themselves Be Constrained?
The findings in Sec. D.2 show that unconstrained optimization can lead to "replication" of an arbitrary image. In our work we default to restricting the number of optimization steps to prevent that false replication. We validate the soundness of that limit empirically, showing in Tab. 1 (second row) that we alleviate the problem. An alternative approach to prevent triggering arbitrary data would be to investigate the embedding space itself, and based on its characteristics, define meaningful constraints on the optimization.
We focus on L2 norms of the embeddings, as in Fig. 5 (right) we observe that the norms stay low for memorized content, while to replicate non-memorized content, the embeddings have to have their norm significantly increased. To get a glimpse at the embedding space, we embed 400k prompts from COCO 2014 train dataset, and note that the distribution of the norms is centered around 250, with standard deviation of around 2. Additionally, we perform gradient optimization of tokens to obtain minimal and maximal possible L2 norm of a text embedding. We find discrete inputs that lower L2 norm down to 220, and inputs that are able to increase the norm to 280.
@@ D.3 Can the Embeddings Themselves Be Constrained?
Appendix E Additional Experiments on Pruning-Based Mitigation
@@
E.1 Hyperparameters
@@
NeMo: We set the memorization score threshold to $\tau_{\text{mem}}=0.428$, corresponding to the mean SSIM score plus one standard deviation, as measured on a holdout set of 50,000 LAION prompts. For the stronger variant of NeMo, reported in Tab. 2, we lowered the threshold to $\tau_{\text{mem}}=0.288$, which corresponds to the mean SSIM score minus one standard deviation. While we follow the original evaluation procedure by individually identifying and disabling neurons for each memorized prompt, we compute the FID and KID metrics by simultaneously deactivating all neurons identified for VM and TM prompts, respectively. This approach provides a more consistent estimate of the pruning’s overall impact on model utility.
Wanda: For Wanda, we follow the experimental setup of Chavhan et al. [3]. Specifically, we use all 500 memorized prompts to identify weights in the second fully connected layer of the cross-attention mechanism. As in Chavhan et al. [3], we select the top $1\%$ of weights with the highest Wanda scores. These weights are then aggregated across the first 10 time steps and pruned to mitigate memorization. Additional results for identifying weights using Wanda per memorized prompt, for 10 and for all 500 memorized prompts, can be seen in Tab. 4. Results for different number of time steps and different values of sparsity can be found in Tab. 5 and Tab. 6, respectively.
E.2 Sensitivity Analysis of Adversarial Embedding Optimization
In Tab. 2, we compare the results of NeMo and Wanda with and without adversarial embedding optimization. Application of adversarial embeddings is denoted by [Uncaptioned image]. Additionally, we repeat the experiments using different numbers of adversarial optimization steps, denoted by Adv. Steps in the table. All optimizations are initialized from the memorized training embedding. Notably, a single optimization step is already sufficient to circumvent the mitigation introduced by NeMo. In the case of Wanda, approximately 25 optimization steps are required before clear replication is triggered.
@@ E.2 Sensitivity Analysis of Adversarial Embedding Optimization
Setting Adv. Steps Memorization Type $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$ $\boldsymbol{\downarrow}$FID $\boldsymbol{\downarrow}$KID
No Mitigation Verbatim $0.90\pm 0.04$ N/A $1.00\pm 0.00$ $0.33\pm 0.01$ 14.44 0.0061
Template $0.17\pm 0.09$ N/A $0.90\pm 0.08$ $0.33\pm 0.02$
NeMo [14] Verbatim $0.33\pm 0.18$ $0.40\pm 0.21$ $0.46\pm 0.13$ $0.34\pm 0.02$ $15.16$ $0.0061$
Template $0.23\pm 0.08$ $0.54\pm 0.28$ $0.55\pm 0.10$ $0.34\pm 0.03$ $18.97$ $0.0048$
Wanda [3] Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $0.34\pm 0.02$ $16.86$ $0.0065$
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $0.34\pm 0.03$ $17.51$ $0.0070$
NeMo + [Uncaptioned image] $1$ Verbatim $0.86\pm 0.07$ $0.94\pm 0.04$ $1.00\pm 0.00$ $0.32\pm 0.01$ $15.16$ $0.0061$
Template $0.28\pm 0.11$ $0.51\pm 0.28$ $0.62\pm 0.21$ $0.33\pm 0.02$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $10$ Verbatim $0.81\pm 0.06$ $0.88\pm 0.05$ $0.99\pm 0.01$ $0.32\pm 0.01$ $15.16$ $0.0061$
Template $0.42\pm 0.13$ $0.21\pm 0.15$ $0.72\pm 0.13$ $0.32\pm 0.02$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $25$ Verbatim $0.88\pm 0.04$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.01$ $15.16$ $0.0061$
Template $0.50\pm 0.12$ $0.20\pm 0.15$ $0.75\pm 0.10$ $0.32\pm 0.02$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $50$ Verbatim $0.91\pm 0.03$ $0.97\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $15.16$ $0.0061$
Template $0.55\pm 0.12$ $0.17\pm 0.12$ $0.79\pm 0.11$ $0.32\pm 0.02$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $100$ Verbatim $0.93\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.32\pm 0.02$ $15.16$ $0.0061$
Template $0.60\pm 0.14$ $0.17\pm 0.12$ $0.86\pm 0.09$ $0.32\pm 0.02$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $150$ Verbatim $0.92\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.32\pm 0.02$ $15.16$ $0.0061$
Template $0.65\pm 0.16$ $0.17\pm 0.12$ $0.93\pm 0.06$ $0.32\pm 0.02$ $18.97$ $0.0048$
NeMo (strong, $\tau_{\text{mem}}=0.288$) + [Uncaptioned image] $50$ Verbatim $0.91\pm 0.03$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $14.92$ $0.0064$
Template $0.55\pm 0.12$ $0.19\pm 0.12$ $0.79\pm 0.10$ $0.32\pm 0.02$ $18.85$ $0.0042$
Wanda + [Uncaptioned image] $1$ Verbatim $0.11\pm 0.05$ $0.11\pm 0.06$ $0.58\pm 0.08$ $0.24\pm 0.04$ $16.86$ $0.0065$
Template $0.11\pm 0.05$ $0.12\pm 0.05$ $0.54\pm 0.09$ $0.31\pm 0.06$ $17.51$ $0.0070$
Wanda + [Uncaptioned image] $10$ Verbatim $0.58\pm 0.11$ $0.64\pm 0.11$ $0.76\pm 0.14$ $0.31\pm 0.02$ $16.86$ $0.0065$
Template $0.12\pm 0.05$ $0.43\pm 0.18$ $0.53\pm 0.16$ $0.31\pm 0.03$ $17.51$ $0.0070$
Wanda + [Uncaptioned image] $25$ Verbatim $0.69\pm 0.07$ $0.77\pm 0.05$ $0.90\pm 0.07$ $0.32\pm 0.02$ $16.86$ $0.0065$
Template $0.12\pm 0.05$ $0.65\pm 0.07$ $0.78\pm 0.10$ $0.32\pm 0.02$ $17.51$ $0.0070$
Wanda + [Uncaptioned image] $50$ Verbatim $0.76\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $16.86$ $0.0065$
Template $0.10\pm 0.06$ $0.73\pm 0.05$ $0.88\pm 0.06$ $0.32\pm 0.02$ $17.51$ $0.0070$
Wanda + [Uncaptioned image] $100$ Verbatim $0.80\pm 0.05$ $0.85\pm 0.04$ $0.98\pm 0.01$ $0.32\pm 0.02$ $16.86$ $0.0065$
Template $0.09\pm 0.06$ $0.75\pm 0.05$ $0.94\pm 0.04$ $0.32\pm 0.02$ $17.51$ $0.0070$
Wanda + [Uncaptioned image] $150$ Verbatim $0.81\pm 0.04$ $0.85\pm 0.04$ $0.99\pm 0.01$ $0.32\pm 0.02$ $16.86$ $0.0065$
Template $0.09\pm 0.06$ $0.76\pm 0.05$ $0.95\pm 0.03$ $0.31\pm 0.02$ $17.51$ $0.0070$
Non-Memorized Images None $0.17\pm 0.05$ N/A $0.35\pm 0.06$ $0.35\pm 0.02$ $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $1$ None $0.17\pm 0.04$ N/A $0.34\pm 0.06$ $0.34\pm 0.02$ $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $10$ None $0.28\pm 0.05$ N/A $0.48\pm 0.06$ $0.32\pm 0.02$ $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $25$ None $0.39\pm 0.06$ N/A $0.58\pm 0.06$ $0.32\pm 0.02$ $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $50$ None $0.48\pm 0.06$ N/A $0.67\pm 0.07$ $0.32\pm 0.02$ $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $100$ None $0.58\pm 0.06$ N/A $0.79\pm 0.07$ $0.32\pm 0.02$ $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $150$ None $0.65\pm 0.06$ N/A $0.88\pm 0.06$ $0.32\pm 0.02$ $14.44$ $0.0061$
E.3 Starting Embedding Optimization from Random Embeddings

### v2

@@ F.1 Can We Make a DM to Output Any Image With Adversarial Embeddings?
F.2 Comparing Behavioral Differences Between Sets
Our findings from Section˜F.1 undermine our adversarial-based memorization identification. In effect, it may seem that our results regarding NeMo and Wanda (Sec.˜3.3) locality in the embedding space (Sec.˜4.1). and locality in the model’s activations (Sec.˜4.2) and weights (Sec.˜4.3) become invalid. Indeed, if we are able to trigger generation of any image, then we should not claim that NeMo and Wanda only conceal the memorization instead of fully removing it, and the findings regarding localization would be false, as the obtained adversarial embeddings ${\bm{y}}_{adv}$ yield little information about how the model (and the embedding space) behaves when faced with memorized data.
@@ F.2 Comparing Behavioral Differences Between Sets
In Fig.˜6 (right) we show how the SSCD score progresses with the optimization. We note that for verbatim memorization (and generated template memorization) we need only a handful of optimization steps to obtain ${\bm{y}}_{adv}$ that reliably triggers generation of the images. For non-memorized data we reach SSCD above $0.7$ after approximately 500 steps. Notably, we require as much as 200 steps to craft ${\bm{y}}_{adv}$ that reliably produces generated (non-memorized) images.
These results show that the optimization process is indeed different for memorized images than for other images. Building on these findings we allow the optimization procedure to only perform $50$ update steps—a value that guarantees generation of memorized images (if present in the model), while preventing False Positives, i.e., generation of non-memorized images. This constraint ensures methodological correctness of our adversarial-based approach, and proves our results in Secs.˜3.3 and 4 are valid.
@@
F.3 Can the Embeddings Themselves Be Constrained?
The findings in Section˜F.2 show that unconstrained optimization can lead to "replication" of an arbitrary image. In our work we default to restricting the number of optimization steps to prevent that false replication. We validate the soundness of that limit empirically, showing in Tab.˜1 (second row) that we alleviate the problem. An alternative approach to prevent triggering arbitrary data would be to investigate the embedding space itself, and based on its characteristics, define meaningful constraints on the optimization.
We focus on L2 norms of the embeddings, as in Fig.˜6 (right) we observe that the norms stay low for memorized content, while to replicate non-memorized content, the embeddings have to have their norm significantly increased. To get a glimpse at the embedding space, we embed 400k prompts from COCO 2014 train dataset, and note that the distribution of the norms is centered around 250, with standard deviation of around 2. Additionally, we perform gradient optimization of tokens to obtain minimal and maximal possible L2 norm of a text embedding. We find discrete inputs that lower L2 norm down to 220, and inputs that are able to increase the norm to 280.
@@ F.3 Can the Embeddings Themselves Be Constrained?
F.4 How do Adversarial Embeddings Relate to Natural Prompts?
In our experiments (Sections˜4.1 and J.1), we initialize the Dori optimization from either randomly sampled embeddings (Fig.˜2), or text embeddings obtained from real, non-memorized text prompts (Fig.˜8). In both cases, we observe that a memorized image can be successfully triggered from any place in the embedding space. Importantly, almost every adversarial embedding stays close to its initialization, be it a random vector or a real text embedding. This finding suggests that adversarial embeddings that cause replication of memorized images do not relate to any semantically meaningful text embeddings, and are instead uniformly scattered in the text embedding space.
Appendix G Concept Unlearning is not Suitable for Memorization Mitigation Research
Concept unlearning techniques aim to permanently suppress the generation of high-level concepts, such as nudity. Although related to the problem of undesired data memorization and replication, we show in this section that existing concept unlearning methods are failing to detect and prevent data replication.
G.1 UnlearnDiffAtk Is Not Capable of Finding Data Replication Triggers
In Section˜3, we introduce our method for searching adversarial embeddings that trigger data replication, even after pruning-based mitigation strategies are applied. While similar strategies have been proposed to critically evaluate concept unlearning techniques, we show that UnlearnDiffAtk (Zhang et al., 2024c), the state-of-the-art adversarial evaluation method for concept unlearning, is not suitable for identifying data replication triggers.
[…]
UnlearnDiffAtk is unsuccessful in breaking pruning-based mitigation methods, i.e., no memorized images are generated from prompt $\bm{p}$$^{\prime}$ after a mitigation is applied. Its failure highlights a core problem with concept removal methods: they focus on discrete text space, instead of exploring the full risk surface, like continuous text embeddings. While for content moderation, such a scope is sufficient, memorization is a privacy risk, which should be evaluated under the worst-case, unconstrained scenario.
G.2 Concept Removal Methods fail To Robustly Mitigate Memorization Against Dori
Beyond using concept unlearning techniques to detect memorization in DMs, which we find to be ineffective in practice, a natural question is whether methods from concept unlearning can be leveraged to address data replication. However, we show that even state-of-the-art approaches are unsuitable for mitigating memorization, underscoring the need for novel methods specifically designed for preventing data replication.
We compare the mitigation capabilities of ESD (Gandikota et al., 2023) and Concept Ablation (Kumari et al., 2023) against our adversarial fine-tuning–based mitigation strategy, with results reported in Tab.˜3. Both concept unlearning methods reduce the extent of data replication, as reflected by the SSCD scores. However, compared to our approach, a larger portion of memorization remains. Moreover, Concept Ablation degrades model utility, as evidenced by the higher FID score.
While these two methods seem to fail to fully remove memorization, they are also quite computational expensive. While ESD could be applied directly, guiding the model away from memorized samples, Concept Ablation requires paraphrasing memorized prompts via ChatGPT to generate a set of anchor prompts. For 7 highly memorized prompts, no valid paraphrase could be generated that did not trigger the memorized image, so these memorized images remained in the model—highlighting a limitation for using Concept Ablation to remove verbatim memorization compared to our method, which directly targets such cases. Since we did not observe strong success in removing memorization when running the methods for their original runtimes, we ran both of them significantly longer until we observed a plateau in their success: ESD was fine‑tuned for 10k steps (10 times longer than in their paper) and Concept Ablation for 12k steps (15 times longer than in their paper). ESD completed in 1 hour and 7 minutes while Concept Ablation required 17 hours 45 minutes due to generating 1000 paraphrased images per prompt (roughly 8 minutes per prompt).
The most striking difference emerges when evaluating the methods with adversarial embeddings, denoted by the [Uncaptioned image] symbol. While our mitigation method remains robust, both ESD and Concept Ablation expose memorized content under adversarial embeddings, suggesting that these methods conceal rather than truly mitigate memorization.
Table 3: Full adversarial fine-tuning is a robust memorization mitigation technique. Contrary to naive application of methods derived from concept removal area, we show that a tailored approach can successfully address memorization in a DM. By resigning from a faulty locality idea, and incorporating diverse set of triggers alongside full weight update we show that robust mitigation is possible.
[…]
ESD (Gandikota et al., 2023) Verbatim $0.24\pm 0.09$ $0.27\pm 0.11$ $0.34\pm 0.02$ 0.11 $13.45$
ESD + [Uncaptioned image] Verbatim $\bm{0.90}\pm 0.04$ $0.97\pm 0.02$ $0.33\pm 0.01$ $\bm{0.98}$
Concept Ablation (Kumari et al., 2023) Verbatim $0.39\pm 0.18$ $0.49\pm 0.25$ $0.34\pm 0.02$ 0.29 $16.03$
Concept Ablation + [Uncaptioned image] Verbatim $\bm{0.91}\pm 0.04$ $0.96\pm 0.02$ $0.32\pm 0.02$ $\bm{0.97}$
Our Mitigation Verbatim $0.15\pm 0.07$ $0.15\pm 0.07$ $33\pm 0.01 0.0$0 $13.61$
[…]
Appendix H Additional Experiments on Pruning-Based Mitigation
@@
We extend our evaluation by diversity, since typically, memorized images are consistently replicated, regardless of the choice of the initial noise $\bm{\epsilon}$. Conversely, for any other input, images generated by a DM are diverse under varying initial noise. We quantify diversity as the average pairwise cosine similarity $\text{D}_{\text{SSCD}}$ between SSCD embeddings of images generated from the same input but different initial noise. Images generated after mitigation should exhibit greater diversity, indicated by lower $\text{D}_{\text{SSCD}}$ values.
H.1 Hyperparameters
@@
NeMo: We set the memorization score threshold to $\tau_{\text{mem}}=0.428$, corresponding to the mean SSIM score plus one standard deviation, as measured on a holdout set of 50,000 LAION prompts. For the stronger variant of NeMo, reported in Table˜4, we lowered the threshold to $\tau_{\text{mem}}=0.288$, which corresponds to the mean SSIM score minus one standard deviation. While we follow the original evaluation procedure by individually identifying and disabling neurons for each memorized prompt, we compute the FID and KID metrics by simultaneously deactivating all neurons identified for VM and TM prompts, respectively. This approach provides a more consistent estimate of the pruning’s overall impact on model utility.
Wanda: For Wanda, we follow the experimental setup of Chavhan et al. (2024). Specifically, we use all 500 memorized prompts to identify weights in the second fully connected layer of the cross-attention mechanism. As in Chavhan et al. (2024), we select the top $1\%$ of weights with the highest Wanda scores. These weights are then aggregated across the first 10 time steps and pruned to mitigate memorization. Additional results for identifying weights using Wanda per memorized prompt, for 10 and for all 500 memorized prompts, can be seen in Table˜6. Results for different number of time steps and different values of sparsity can be found in Table˜7 and Table˜8, respectively.
H.2 Sensitivity Analysis of Adversarial Embedding Optimization
In Table˜4, we compare the results of NeMo and Wanda with and without adversarial embedding optimization. Application of adversarial embeddings is denoted by [Uncaptioned image]. Additionally, we repeat the experiments using different numbers of adversarial optimization steps, denoted by Adv. Steps in the table. All optimizations are initialized from the memorized training embedding. Notably, a single optimization step is already sufficient to circumvent the mitigation introduced by NeMo. In the case of Wanda, approximately 25 optimization steps are required before clear replication is triggered.
@@ H.2 Sensitivity Analysis of Adversarial Embedding Optimization
[…]
No Mitigation Verbatim $0.90\pm 0.04$ N/A $1.00\pm 0.00$ $0.33\pm 0.01$ $1.00$ 14.44 0.0061
Template $0.17\pm 0.09$ N/A $0.90\pm 0.08$ $0.33\pm 0.02$ $1.00$
NeMo (Hintersdorf et al., 2024) Verbatim $0.33\pm 0.18$ $0.40\pm 0.21$ $0.46\pm 0.13$ $0.34\pm 0.02$ $0.20$ $15.16$ $0.0061$
Template $0.23\pm 0.08$ $4\pm 0.28 0.55±0$$5\pm 0.10$ $0.34\pm 0.03$ $0.15$ $18.97$ $0.0048$
Wanda (Chavhan et al., 2024) Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $34\pm 0.02 0.0$00.00 $16.86$ $0.0065$
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $34\pm 0.03 0.0$00.00 $17.51$ $0.0070$
NeMo + [Uncaptioned image] 11 Verbatim $0.86\pm 0.07$ $0.94\pm 0.04$ $1.00\pm 0.00$ $0.32\pm 0.01$ $0.73$ $15.16$ $0.0061$
Template $0.28\pm 0.11$ $0.51\pm 0.28$ $0.62\pm 0.21$ $0.33\pm 0.02$ $0.21$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $10$ Verbatim $1\pm 0.06 0.88±0$$8\pm 0.05$ $0.99\pm 0.01$ $0.32\pm 0.01$ $0.79$ $15.16$ $0.0061$
Template $0.42\pm 0.13$ $0.21\pm 0.15$ $0.72\pm 0.13$ $0.32\pm 0.02$ $0.14$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $25$ Verbatim $0.88\pm 0.04$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.01$ $0.94$ $15.16$ $0.0061$
Template $0.50\pm 0.12$ $0.20\pm 0.15$ $0.75\pm 0.10$ $0.32\pm 0.02$ $0.13$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $50$ Verbatim $0.91\pm 0.03$ $0.97\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $0.99$ $15.16$ $0.0061$
Template $0.55\pm 0.12$ $0.17\pm 0.12$ $0.79\pm 0.11$ $0.32\pm 0.02$ $0.12$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $100$ Verbatim $0.93\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.32\pm 0.02$ $1.00$ $15.16$ $0.0061$
Template $0.60\pm 0.14$ $0.17\pm 0.12$ $0.86\pm 0.09$ $0.32\pm 0.02$ $0.10$ $18.97$ $0.0048$
NeMo + [Uncaptioned image] $150$ Verbatim $0.92\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.32\pm 0.02$ $0.99$ $15.16$ $0.0061$
Template $0.65\pm 0.16$ $0.17\pm 0.12$ $0.93\pm 0.06$ $32\pm 0.02 0.0$80.08 $18.97$ $0.0048$
NeMo (strong, $\tau_{\text{mem}}=0.288$) + [Uncaptioned image] $50$ Verbatim $0.91\pm 0.03$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $0.88$ $14.92$ $0.0064$
Template $0.55\pm 0.12$ $0.19\pm 0.12$ $0.79\pm 0.10$ $0.32\pm 0.02$ $0.15$ $18.85$ $0.0042$
Wanda + [Uncaptioned image] 11 Verbatim $1\pm 0.05 0.11±0$$1\pm 0.06$ $0.58\pm 0.08$ $24\pm 0.04 0.0$20.02 $16.86$ $0.0065$
Template $0.19\pm 0.07$ $0.16\pm 0.07$ $0.51\pm 0.11$ $32\pm 0.03 0.0$00.00 $17.51$ $0.0070$
Wanda + [Uncaptioned image] $10$ Verbatim $0.58\pm 0.11$ $0.64\pm 0.11$ $0.76\pm 0.14$ $0.31\pm 0.02$ $0.18$ $16.86$ $0.0065$
Template $0.37\pm 0.10$ $0.17\pm 0.09$ $0.61\pm 0.12$ $32\pm 0.03 0.0$20.02 $17.51$ $0.0070$
Wanda + [Uncaptioned image] $25$ Verbatim $0.69\pm 0.07$ $0.77\pm 0.05$ $0.90\pm 0.07$ $0.32\pm 0.02$ $0.49$ $16.86$ $0.0065$
Template $0.45\pm 0.11$ $0.17\pm 0.10$ $0.70\pm 0.09$ $32\pm 0.03 0.0$70.07 $17.51$ $0.0070$
Wanda + [Uncaptioned image] $50$ Verbatim $0.76\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $0.73$ $16.86$ $0.0065$
Template $0.51\pm 0.11$ $0.16\pm 0.09$ $0.75\pm 0.08$ $0.32\pm 0.02$ $0.16$ $17.51$ $0.0070$
Wanda + [Uncaptioned image] $100$ Verbatim $0.80\pm 0.05$ $0.85\pm 0.04$ $0.98\pm 0.01$ $0.32\pm 0.02$ $0.86$ $16.86$ $0.0065$
Template $0.53\pm 0.12$ $0.15\pm 0.08$ $0.78\pm 0.08$ $0.32\pm 0.02$ $0.26$ $17.51$ $0.0070$
Wanda + [Uncaptioned image] $150$ Verbatim $0.81\pm 0.04$ $0.85\pm 0.04$ $0.99\pm 0.01$ $0.32\pm 0.02$ $0.91$ $16.86$ $0.0065$
Template $0.54\pm 0.14$ $0.15\pm 0.09$ $0.82\pm 0.08$ $\pm 0.02 0.320.32$ $17.51$ $0.0070$
Non-Memorized Images None $0.17\pm 0.05$ N/A $0.35\pm 0.06$ $35\pm 0.02 0.0$00.00 $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] 11 None $0.17\pm 0.04$ N/A $0.34\pm 0.06$ $34\pm 0.02 0.0$00.00 $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $10$ None $0.28\pm 0.05$ N/A $0.48\pm 0.06$ $32\pm 0.02 0.0$00.00 $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $25$ None $0.39\pm 0.06$ N/A $0.58\pm 0.06$ $32\pm 0.02 0.0$00.00 $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $50$ None $0.48\pm 0.06$ N/A $0.67\pm 0.07$ $32\pm 0.02 0.0$20.02 $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $100$ None $0.58\pm 0.06$ N/A $0.79\pm 0.07$ $32\pm 0.02 0.0$80.08 $14.44$ $0.0061$
Non-Memorized Images + [Uncaptioned image] $150$ None $0.65\pm 0.06$ N/A $0.88\pm 0.06$ $0.32\pm 0.02$ $0.25$ $14.44$ $0.0061$
H.3 Starting Embedding Optimization from Random Embeddings